Exercise 21.14 Constants Part A A charge of -0.500 C exerts an upward 0.220-N force on an unknown charge 0.320 m directly below it. What is the unknown charge (magnitude and sign)? e- Submit Request Answer Part B What is the magnitude of the force that the unknown charge exerts on the charge of magnitude-0.500 μ C ? Submit Request Answer
Answer #1

a)

F = k q1 q2 / r^2

0.22 = 9*10^9* 0.5*10^-6* q2 / 0.32^2

q2 = - 5*10^-6 C

b)

F = 9*10^9* 5*10^-6* 0.5*10^-6 / 0.32^2

F = 0.22 N

=========